OpenStudy (anonymous):

A dripping faucet wastes a cup of water if it drips for four minutes. The amount of water wasted varies directly with the amount of time the faucet drips. What is the constant of variation? a. 4 b. 1/2 c. 2 d. 1/4

OpenStudy (anonymous):

amount of water wasted varies directly with the amount of time w varies directly with t w = kt where k is constant of proportionality 1 = k × 4 k = 1/4 cup per minute ----

OpenStudy (anonymous):

So D